# Intrinsic Saliency Guided Trunk-Collateral Network for Unsupervised Video Object Segmentation

### v1

@@ Abstract
Recent unsupervised video object segmentation (UVOS) methods predominantly adopt the motion-appearance paradigm, which leverages optical flow for motion cues and RGB image for appearance information to facilitate mask prediction. Mainstream motion-appearance approaches use either the two-encoder structure to separately encode motion and appearance features, or the single-encoder structure for joint encoding. However, these methods fail to properly balance the motion-appearance relationship. Consequently, even with complex fusion modules for motion-appearance integration, the extracted suboptimal features degrade the models’ overall performance. Moreover, the quality of optical flow varies across scenarios, making it insufficient to rely solely on optical flow to achieve high-quality segmentation results. To address these challenges, we propose the Intrinsic Saliency guided Trunk-Collateral Network (ISTC-Net), which better balances the motion-appearance relationship and incorporates model’s intrinsic saliency information to enhance segmentation performance. Specifically, considering that optical flow maps are derived from RGB images, they share both commonalities and differences. We propose a novel Trunk-Collateral structure. The shared trunk backbone captures the motion-appearance commonality, while the collateral branch learns the uniqueness of motion features. Furthermore, an Intrinsic Saliency guided Refinement Module (ISRM) is devised to efficiently leverage the model’s intrinsic saliency information to refine high-level features, and provide pixel-level guidance for motion-appearance fusion, thereby enhancing performance without additional input. Experimental results show that ISTC-Net achieved state-of-the-art performance on three UVOS datasets (89.2% $\mathcal{J}\&\mathcal{F}$ on DAVIS-16, 76% $\mathcal{J}$ on YouTube-Objects, 86.4% $\mathcal{J}$ on FBMS) and four standard video salient object detection (VSOD) benchmarks with the notable increase, demonstrating its effectiveness and superiority over previous methods.
@@
(a) Two-Encoder Architecture
(b) Single-Encoder Architecture
@@
Figure 1: Demonstration of typical motion-appearance UVOS paradigms and our trunk-collateral architecture.
@@ I Introduction
Despite recent advances in motion-appearance techniques, several critical challenges still persist. First of all, the general architectures of motion-appearance approaches do not accommodate the relationship between motion and appearance features. Recent motion-appearance methods can be categorized into two types: two-encoder architecture and single-encoder architecture. As depicted in Fig. 1a, the two-encoder architecture[27, 32, 33, 35, 40, 41, 49] employs distinct encoders to separately process RGB images and optical flow maps, yet overlooks the inherent motion-appearance correlations, leading to modality misalignment. In contrast, the single-encoder architecture[37, 42, 43], as shown in Fig. 1b, encodes both features concurrently to reduce computational burden, but fails to capture the motion-appearance distinctions adequately. Neither of these two architectures can well enough balance motion-appearance relationship. Accordingly, previous methods often adopted complex interaction or fusion modules, either at the encoding stage or at the decoding stage, to compensate for the inferior feature extraction. Complicated fusion schemes bring about heavy model design, yet can not fundamentally resolve the problem, eventually leading to unsatisfactory performance.
@@
Figure 2: Examples of optical flow maps in scenarios with partial estimation failure or suboptimal performance. Cases 1–4 respectively represent stationary objects, motion blur, co-moving background, and misidentification of background objects.
Secondly, relying solely on optical flow as auxiliary information is inadequate. Optical flow may fail to capture motion cues in scenarios involving stationary objects, motion blur, or co-moving background elements, as illustrated in Fig. 2. Furthermore, in the absence of instructive information, conventional fusion modules struggle to handle low-quality motion features. To address these limitations, additional information is essential. Saliency maps do not depend on motion cues to detect primary objects, which offer a natural complement in circumstances where optical flow fails. For example, saliency maps can accurately identify stationary foreground objects and filter out co-moving indistinctive elements. However, directly employing external saliency features, such as pre-generated saliency maps, is rather cumbersome. And simply integrating a new feature, via a three-stream structure for instance, may introduce significant computational and memory overhead. Therefore, we argue that exploring a more efficient pattern to incorporate saliency features is worth further investigation.
In this study, we propose a novel intrinsic saliency guided trunk-collateral network (ISTC-Net) which effectively addresses the aforementioned challenges. To address the first challenge, we introduce a brand-new trunk-collateral architecture tailored for motion-appearance UVOS methods. This architecture, as shown in Fig. 1c, leverages a shared backbone to jointly encode motion and appearance features, which captures their shared characteristics; complemented by a set of collateral branches to extract the motion-specific features that are distinct from those shared with appearance features. This innovative pattern simultaneously accounts for both motion-appearance similarity and divergence, eliminating the need for complex fusion modules. Theoretically, the parameter size of the collateral branch should remain relatively small to avoid negative impacts on the learning capacity of the trunk section. At the same time, it must sufficiently capture the motion uniqueness. Low-rank adaptation (LoRA)[50] is a low-rank and lightweight learning structure that can effectively captures additional information without interfering with the main structure. Its low-rank design also enables significant parameter reduction. Therefore, we implement LoRA in the design of the collateral branches, which realizes an optimal balance between computational efficiency and accuracy.
@@
1.We propose an intrinsic saliency guided trunk-collateral network (ISTC-Net) which simultaneously adopts motion, appearance and saliency features for UVOS, and possesses better potential to manage the motion-appearance relationship and effectively incorporate saliency features.
@@
3.Extensive experiments demonstrate the superiority of our ISTC-Net. ISTC-Net achieved state-of-the-art (SOTA) performance across all three widely used UVOS datasets (89.2% $\mathcal{J}\&\mathcal{F}$ on DAVIS-16[1], 76% $\mathcal{J}$ on YouTube-Objects[2], and 86.4% $\mathcal{J}$ on FBMS[3]) and four standard video salient object detection (VSOD) benchmarks.
@@ II-A Incremental Structure
Incremental structure is also widely applied in parameter-efficient fine-tuning. Large language models (LLM) [75, 76, 77, 78, 79] have become foundation models for various downstream applications, due to their striking performance in language understanding and processing. However, LLMs often have massive parameters, requiring substantial computational resources for full fine-tuning, which underscores the significance of PEFT. Some PEFT algorithms reduce the cost of fine-tuning by adding trainable parameters to the frozen baselines. For example, adapter-based methods [80, 81, 82, 83, 84, 85] insert trainable adapters among layers in the Transformer block, allowing the model to adjust to new tasks. LoRA-based approaches [50, 88, 89, 90, 91, 92] add trainable modules comprising two low-rank matrices aside the models’ fully connected layers, which empowers the model to accommodate to new assignments while remarkably reduces the computational burden.
Our work innovatively applies the incremental structure to UVOS. By devising the trunk-collateral structure based on the intrinsic relationship between optical flow and image features, our ISTC-Net accounts for both motion-appearance commonality and differences, thus considerably enhancing the model performance.
@@ II-B Unsupervised Video Object Segmentation
Our ISTC-Net differs from traditional motion-appearance UVOS approaches by designing a novel trunk-collateral structure which better corresponds to the intrinsic relationship between motion and appearance features, thus eliminating the need for complex fusion mechanisms. Furthermore, we exploit the model’s inherent saliency to optimize high-level representations and motion-appearance combination, which further elevates the model effects.
@@ III Method
Figure 3: Overall pipeline of ISTC-Net, which consists of a trunck-collateral encoder, an intrinsic saliency guided refinement module (ISRM) and a multi-level decoder. Given an input image and its corresponding optical flow, the trunk-collateral encoder captures the shared attributes of motion and appearance features through the common trunk section, while encoding distinctive motion characteristics via the collateral branches alongside the trunk section. ISRM optimizes feature representations and enhances motion-appearance integration using the intrinsic saliency feature generated in the first decoding round. The ultimate segmentation result is output in the second decoding round.
@@ III-A Overall Pipeline
In this section, we first provide a comprehensive overview of our model pipeline, which is illustrated in Fig. 3. Specifically, the proposed intrinsic saliency guided trunk-collateral network (ISTC-Net) is an end-to-end UVOS framework based on the encoder-decoder architecture. Given an input video frame $I\in\mathbb{R}^{H\times W\times 3}$ and its corresponding optical flow map $O\in\mathbb{R}^{H\times W\times 3}$, ISTC-Net processes $I$ and $O$ to generate a segmentation mask $M_{\text{pred}}\in\mathbb{R}^{H\times W\times 1}$.
@@ III-B Trunk-Collateral Paradigm
As illustrated in Fig. 1c and Fig. 3, we propose a novel trunk-collateral paradigm for motion-appearance UVOS, formulated to effectively capture both the similarities and divergences between motion and appearance features. The trunk component, denoted as $P_{\text{trunk}}$, serves as a shared backbone responsible for encoding both motion and appearance features, thereby facilitating the extraction of their common characteristics. Complementing this, the collateral branch $P_{\text{collateral}}$ is specifically assigned for processing motion features, isolating the unique characteristics of motion. These distinct motion features are then integrated with the trunk-encoded motion characteristics to generate the final motion representation. Following prior works [37, 41, 46], we adopt Segformer[56] as the backbone, , which constitutes the trunk component $P_{\text{trunk}}$. To align with the Transformer-based [57] architecture, the collateral branch $P_{\text{collateral}}$ incorporated into every dense layer within the Transformer blocks, including the feed-forward network (FFN) and multi-head self-attention (MHSA). Fig. 4 provides a detailed visualization of the trunk-collateral paradigm applied within the Transformer-like structure. Specifically, Fig. 4 (a) and Fig. 4 (b) illustrate the implementation of the trunk-collateral paradigm within the FFN and MHSA, respectively.
@@
Specifically, a linear layer is typically applied to $I_{i}\in\mathbb{R}^{H_{i}W_{i}\times C_{i}}$ and $O_{i}\in\mathbb{R}^{H_{i}W_{i}\times C_{i}}$ to generate their key, query, and value features in each Transformer block. In addition, an auxiliary module is incorporated alongside this linear layer as the collateral branch $P_{\text{collateral}}$ exclusively for processing $O_{i}$. Taking key generation as an example, the encoding of key features $K^{i}_{I}$ and $K^{i}_{O}$ for $I_{i}$ and $O_{i}$ respectively, is performed as expressed in (4).
@@
As demonstrated in (6), this process encompasses the scenario of directly integrating $P_{\text{collateral}}$ to FFN, while simultaneously providing greater flexibility for adaptive learning.
@@ III-C Intrinsic Saliency Guided Refinement Module
The similarity maps $W_{o}\in\mathbb{R}^{\frac{H}{32}\times\frac{W}{32}}$ and $W_{i}\in\mathbb{R}^{\frac{H}{32}\times\frac{W}{32}}$ are derived by computing the pixel-wise cosine similarity between $O_{4\_\text{key}}$, $I_{4\_\text{key}}$, and $S^{\prime}_{\text{key}}$. Since $S^{\prime}$ provides relatively reliable guidance, the map $W_{o}$ and $W_{i}$ represent the pixel-wise confidence of the feature maps $O_{4}$ and $I_{4}$. Subsequently, $W_{o}$ and $W_{i}$ are normalized and applied as weights to $O_{4}$ and $I_{4}$ for the summation, leading to the final combined feature $F_{4o}$.
@@ IV-A1 Traning and Inference
The training process of ISTC-Net comprises two phases. First, we pre-train the model on the YouTube-VOS dataset [4] for 50 epochs to enhance its generalization capability. Subsequently, the model is fine-tuned on specific downstream datasets, such as DAVIS-16 [1] or FBMS [3], to evaluate its performance on the UVOS task. Furthermore, to assess the model’s effectiveness on the VSOD task, we fine-tune the pre-trained model on a mixed dataset comprising DAVIS-16 and DAVSOD [6] like the preceding works, or directly adopt the fine-tuned UVOS model for testing. We employ the AdamW optimizer [61]. The learning rate is set to $6\times 10^{-5}$ during the pre-training phase, and adjusted to $1\times 10^{-4}$ with a learning rate decay strategy in the fine-tuning phase. To enhance training robustness, data augmentation techniques, such as horizontal flipping, are applied. The model is trained on two NVIDIA 3090 GPUs, with a mini-batch size of 16 during pre-training and 8 during fine-tuning.
@@ IV-A2 Evaluation Metrics
For the VSOD task, we also use standard evaluation metrics to assess the performance of ISTC-Net, including the mean absolute error ($MAE$), maximum F-measure ($F_{m}$), maximum enhanced-alignment measure ($E_{m}$) and structure-measure ($S_{m}$). To be more specific, $MAE$ computes the average absolute error on a per-pixel basis; $F_{m}$ integrates precision and recall under the optimal threshold; $E_{m}$ accounts for both pixel-level and image-level discrepancies, and $S_{m}$ quantifies structural similarity from both region-aware and object-aware perspectives.

### v2

# Saliency-Motion Guided Trunk-Collateral Network for Unsupervised Video Object Segmentation

@@ Abstract
Recent unsupervised video object segmentation (UVOS) methods predominantly adopt the motion-appearance paradigm, which leverages optical flow for motion cues and RGB image for appearance information to facilitate mask prediction. Mainstream motion-appearance approaches use either the bi-encoder structure to separately encode motion and appearance features, or the uni-encoder structure for joint encoding. However, these methods fail to properly balance the motion-appearance relationship. Consequently, even with complex fusion modules for motion-appearance integration, the extracted suboptimal features degrade the models’ overall performance. Moreover, the quality of optical flow varies across scenarios, making it insufficient to rely solely on optical flow to achieve high-quality segmentation results. To address these challenges, we propose the Saliency-Mmotion guided Trunk-Collateral Network (SMTC-Net), which better balances the motion-appearance relationship and incorporates model’s intrinsic saliency information to enhance segmentation performance. Specifically, considering that optical flow maps are derived from RGB images, they share both commonalities and differences.Accordingly, we propose a novel Trunk-Collateral structure for motion-appearance UVOS. The shared trunk backbone captures the motion-appearance commonality, while the collateral branch learns the uniqueness of motion features. Furthermore, an Intrinsic Saliency guided Refinement Module (ISRM) is devised to efficiently leverage the model’s intrinsic saliency information to refine high-level features, and provide pixel-level guidance for motion-appearance fusion, thereby enhancing performance without additional input. Experimental results show that SMTC-Net achieved state-of-the-art performance on three UVOS datasets (89.2% $\mathcal{J}\&\mathcal{F}$ on DAVIS-16, 76% $\mathcal{J}$ on YouTube-Objects, 86.4% $\mathcal{J}$ on FBMS) and four standard video salient object detection (VSOD) benchmarks with the notable increase, demonstrating its effectiveness and superiority over previous methods.
@@
(a) Bi-Encoder Architecture
(b) Uni-Encoder Architecture
@@
Figure 1: Illustration of typical motion-appearance UVOS paradigms and the proposed trunk-collateral architecture. (a) Bi-Encoder architecture with separate encoders for images and optical flows. (b) Uni-Encoder architecture with a shared encoder for images and optical flows. (c) Our Trunk-Collateral architecture which consists of both shared and specific parameters for images and optical flows.
@@ I Introduction
Despite recent advances in motion-appearance techniques, several critical challenges still persist. First of all, the general architectures of motion-appearance approaches do not accommodate the relationship between motion and appearance features. Recent motion-appearance methods can be categorized into two types: bi-encoder architecture and uni-encoder architecture. As depicted in Fig. 1a, the bi-encoder architecture[27, 32, 33, 35, 40, 41, 49] employs distinct encoders to separately process RGB images and optical flow maps, yet overlooks the inherent motion-appearance correlations, leading to modality misalignment. In contrast, the uni-encoder architecture[37, 42, 43], as shown in Fig. 1b, encodes both features concurrently to reduce computational burden, but fails to capture the motion-appearance distinctions adequately. Neither of these two architectures can well enough balance motion-appearance relationship. Accordingly, previous methods often adopted complex interaction or fusion modules, either at the encoding stage or at the decoding stage, to compensate for the inferior feature extraction. Complicated fusion schemes bring about heavy model design, yet can not fundamentally resolve the problem, eventually leading to unsatisfactory performance.
@@
Figure 2: Examples of optical flow maps in scenarios with partial estimation failure or suboptimal performance. Case 1–4 respectively represent stationary objects, motion blur, co-moving background, and misidentification of background objects.
Secondly, relying solely on optical flow as auxiliary information is inadequate. Optical flow may fail to capture motion cues in scenarios involving stationary objects, motion blur, or co-moving background elements, as illustrated in Fig. 2. Furthermore, in the absence of instructive information, conventional fusion modules struggle to handle low-quality motion features. To address these limitations, additional information is essential. Saliency maps do not depend on motion cues to detect primary objects, which offer a natural complement in circumstances where optical flow fails. For example, saliency maps can accurately identify stationary foreground objects and filter out co-moving indistinctive elements. However, directly employing external saliency features, such as pre-generated saliency maps, is rather cumbersome. And simply integrating a new feature, like via a three-stream structure for instance, may introduce significant computational and memory overhead. Therefore, we argue that exploring a more efficient pattern to incorporate saliency features is worth further investigation.
In this study, we propose a novel saliency-motion guided trunk-collateral network (SMTC-Net) which effectively addresses the aforementioned challenges. To address the first challenge, we introduce a brand-new trunk-collateral architecture tailored for motion-appearance UVOS methods. This architecture, as shown in Fig. 1c, leverages a shared backbone to jointly encode motion and appearance features, which captures their shared characteristics; complemented by a set of collateral branches to extract the motion-specific features that are distinct from those shared with appearance features. This innovative pattern simultaneously accounts for both motion-appearance similarity and divergence, eliminating the need for complex fusion modules. Theoretically, the parameter size of the collateral branch should remain relatively small to avoid negative impacts on the learning capacity of the trunk section. At the same time, it must sufficiently capture the motion uniqueness. Low-rank adaptation (LoRA)[50] is a low-rank and lightweight learning structure that can effectively captures additional information. The low-rank design also enables significant parameter reduction which avoids interfering with learning capacity of the main structure. Therefore, we implement LoRA in the design of the collateral branches, which realizes an optimal balance between computational efficiency and accuracy.
@@
1.We propose an saliency-motion guided trunk-collateral network (SMTC-Net) which simultaneously adopts motion, appearance and saliency features for UVOS, and possesses better potential to manage the motion-appearance relationship while effectively incorporating saliency features.
@@
3.Extensive experiments demonstrate the superiority of our SMTC-Net. SMTC-Net achieved state-of-the-art (SOTA) performance across all three widely used UVOS datasets (89.2% $\mathcal{J}\&\mathcal{F}$ on DAVIS-16[1], 76% $\mathcal{J}$ on YouTube-Objects[2], and 86.4% $\mathcal{J}$ on FBMS[3]) and four standard video salient object detection (VSOD) benchmarks.
@@ II-A Incremental Structure
Incremental structure is also widely applied in parameter-efficient fine-tuning. Large language models (LLM) [75, 76, 77, 78, 79] have become foundation models for various downstream applications, due to their striking performance in language understanding and processing. However, LLMs often have massive parameters, requiring substantial computational resources for full fine-tuning, which underscores the significance of PEFT. Some PEFT algorithms reduce the cost of fine-tuning by adding trainable parameters to the frozen baselines. For example, adapter-based methods [80, 81, 82, 83, 84, 85] insert trainable adapters among layers in the Transformer block, allowing the model to adjust to new tasks. LoRA-based approaches [50, 88, 89, 90, 91] add trainable modules comprising two low-rank matrices aside the models’ fully connected layers, which empowers the model to accommodate to new assignments while remarkably reduces the computational burden.
Our work innovatively applies the incremental structure to UVOS. By devising the trunk-collateral structure based on the intrinsic relationship between optical flow and image features, our SMTC-Net accounts for both motion-appearance commonality and differences, thus considerably enhancing the model performance.
@@ II-B Unsupervised Video Object Segmentation
Our SMTC-Net differs from traditional motion-appearance UVOS approaches by designing a novel trunk-collateral structure which better corresponds to the intrinsic relationship between motion and appearance features, thus eliminating the need for complex fusion mechanisms. Furthermore, we exploit the model’s inherent saliency to optimize high-level representations and motion-appearance combination, which further elevates the model effects.
@@ III Method
Figure 3: Overall pipeline of SMTC-Net, which consists of a trunck-collateral encoder, an intrinsic saliency guided refinement module (ISRM) and a multi-level decoder. Given an input image and its corresponding optical flow, the trunk-collateral encoder captures the shared attributes of motion and appearance features through the common trunk section, while encoding distinctive motion characteristics via the collateral branches alongside the trunk section. ISRM optimizes feature representations and enhances motion-appearance integration using the intrinsic saliency feature generated in the first decoding round. The ultimate segmentation result is output in the second decoding round.
@@ III-A Overall Pipeline
In this section, we first provide a comprehensive overview of our model pipeline, which is illustrated in Fig. 3. Specifically, the proposed saliency-motion guided trunk-collateral network (SMTC-Net) is an end-to-end UVOS framework based on the encoder-decoder architecture. Given an input video frame $I\in\mathbb{R}^{H\times W\times 3}$ and its corresponding optical flow map $O\in\mathbb{R}^{H\times W\times 3}$, SMTC-Net processes $I$ and $O$ to generate a segmentation mask $M_{\text{pred}}\in\mathbb{R}^{H\times W\times 1}$.
@@ III-B Trunk-Collateral Paradigm
As illustrated in Fig. 1c and Fig. 3, we propose a novel trunk-collateral paradigm for motion-appearance UVOS, formulated to effectively capture both the similarities and divergences between motion and appearance features. The trunk component, denoted as $P_{\text{trunk}}$, serves as a shared backbone responsible for encoding both motion and appearance features, thereby facilitating the extraction of their common characteristics. Complementing this, the collateral branch $P_{\text{collateral}}$ is specifically assigned for processing motion features, isolating the unique characteristics of motion. These distinct motion features are then integrated with the trunk-encoded motion characteristics to generate the final motion representation. Following prior works [37, 41, 46], we adopt Segformer[56] as the backbone, which constitutes the trunk component $P_{\text{trunk}}$. To align with the Transformer-based [57] architecture, the collateral branch $P_{\text{collateral}}$ is incorporated into every dense layer within the Transformer blocks, including the feed-forward network (FFN) and multi-head self-attention (MHSA). Fig. 4 provides a detailed visualization of the trunk-collateral paradigm applied within the Transformer-like structure. Specifically, Fig. 4 (a) and Fig. 4 (b) illustrate the implementation of the trunk-collateral paradigm within the FFN and MHSA, respectively.
@@
Specifically, a linear layer is typically applied to $I_{i}\in\mathbb{R}^{H_{i}W_{i}\times C_{i}}$ and $O_{i}\in\mathbb{R}^{H_{i}W_{i}\times C_{i}}$ to generate their key, query, and value features during MHSA in each Transformer block. In addition, the collateral branches $P_{\text{collateral}}$ are incorporated alongside these linear layers exclusively for processing $O_{i}$. Taking key generation as an example, the encoding of key features $K^{i}_{I}$ and $K^{i}_{O}$ for $I_{i}$ and $O_{i}$ respectively, is performed as expressed in (4).
@@
As demonstrated in (6), this process encompasses the scenario of directly integrating $P_{\text{collateral}}$ to FFN, while simultaneously providing more flexibility for adaptive learning.
@@ III-C Intrinsic Saliency Guided Refinement Module
The similarity maps $W_{o}\in\mathbb{R}^{\frac{H}{32}\times\frac{W}{32}}$ and $W_{i}\in\mathbb{R}^{\frac{H}{32}\times\frac{W}{32}}$ are derived by computing the pixel-wise cosine similarity between $O_{4\_\text{key}}$, $I_{4\_\text{key}}$, and $S^{\prime}_{\text{key}}$. Since $S^{\prime}$ provides relatively reliable guidance, the map $W_{o}$ and $W_{i}$ represent the pixel-wise confidence of the feature maps $O_{4}$ and $I_{4}$. Subsequently, $W_{o}$ and $W_{i}$ are normalized and applied as the summation weights to $O_{4}$ and $I_{4}$, leading to the final combined feature $F_{4o}$.
@@ IV-A1 Traning and Inference
The training process of SMTC-Net comprises two phases. First, we pre-train the model on the YouTube-VOS dataset [4] for 50 epochs to enhance its generalization capability. Subsequently, the model is fine-tuned on specific downstream datasets, such as DAVIS-16 [1] or FBMS [3], to evaluate its performance on the UVOS task. Furthermore, to assess the model’s effectiveness on the VSOD task, we fine-tune the pre-trained model on a mixed dataset comprising DAVIS-16 and DAVSOD [6] like the preceding works, or directly adopt the fine-tuned UVOS model for testing. We employ the AdamW optimizer [61]. The learning rate is fixed to $6\times 10^{-5}$ during the pre-training phase, and adjusted to $1\times 10^{-4}$ with a learning rate decay strategy in the fine-tuning phase. To enhance training robustness, data augmentation techniques, such as horizontal flipping, are applied. The model is trained on two NVIDIA 3090 GPUs, with a mini-batch size of 16 during pre-training and 8 during fine-tuning.
@@ IV-A2 Evaluation Metrics
For the VSOD task, we also use standard evaluation metrics to assess the performance of SMTC-Net, including the mean absolute error ($MAE$), maximum F-measure ($F_{m}$), maximum enhanced-alignment measure ($E_{m}$) and structure-measure ($S_{m}$). To be more specific, $MAE$ computes the average absolute error on a per-pixel basis; $F_{m}$ integrates precision and recall under the optimal threshold; $E_{m}$ accounts for both pixel-level and image-level discrepancies, and $S_{m}$ quantifies structural similarity from both region-aware and object-aware perspectives.